OpenStudy (anonymous):

Solve the equation: −4(x + 10) − 6 = −3(x − 2

OpenStudy (anonymous):

Solve for x: -6-4 (x+10) = -3 (x-20) Distribute -4 over x+10. -4 (x+10) = -40-4 x: -40-4 x-6 = -3 (x-20) Group like terms in -4 x-40-6. Grouping like terms, -4 x-40-6 = -4 x+(-6-40): -4 x+(-6-40) = -3 (x-20) Evaluate -6-40. -6-40 = -46: -46-4 x = -3 (x-20) Factor the polynomial, -4 x-46. Factor -2 from the polynomial -46-4 x: -2 (2 x+23) = -3 (x-20) Move everything to the left hand side. Add 3 (x-20) to both sides of -2 (2 x+23) = -3 (x-20): 3 (x-20)-2 (2 x+23) = 3 (x-20)-3 (x-20) Look for two terms that sum to zero. 3 (x-20)-3 (x-20) = 0: 3 (x-20)-2 (2 x+23) = 0 Distribute -2 over 2 x+23. -2 (2 x+23) = -46-4 x: -46-4 x+3 (x-20) = 0 Distribute 3 over x-20. 3 (x-20) = 3 x-60: 3 x-60-4 x-46 = 0 Group like terms in 3 x-4 x-60-46. Grouping like terms, 3 x-4 x-60-46 = (3 x-4 x)+(-46-60): (3 x-4 x)+(-46-60) = 0 Combine like terms in 3 x-4 x. 3 x-4 x = -x: -x+(-46-60) = 0 Evaluate -46-60. -46-60 = -106: -106-x = 0 Isolate terms with x to the left hand side. Add 106 to both sides: (106-106)-x = 106 Look for two terms that sum to zero. 106-106 = 0: -x = 106 Solve for x. Multiply both sides of -x = 106 by -1: -x/(-1) = -106 Any nonzero number divided by itself is one. (-1)/(-1) = 1: Answer: | | x = -106 If you liked my repsonse, please give me the best response!

OpenStudy (jonathanvaldes98):

Simplifying -4(x + 10) + -6 = -3(x + -2) Reorder the terms: -4(10 + x) + -6 = -3(x + -2) (10 * -4 + x * -4) + -6 = -3(x + -2) (-40 + -4x) + -6 = -3(x + -2) Reorder the terms: -40 + -6 + -4x = -3(x + -2) Combine like terms: -40 + -6 = -46 -46 + -4x = -3(x + -2) Reorder the terms: -46 + -4x = -3(-2 + x) -46 + -4x = (-2 * -3 + x * -3) -46 + -4x = (6 + -3x) Solving -46 + -4x = 6 + -3x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '3x' to each side of the equation. -46 + -4x + 3x = 6 + -3x + 3x Combine like terms: -4x + 3x = -1x -46 + -1x = 6 + -3x + 3x Combine like terms: -3x + 3x = 0 -46 + -1x = 6 + 0 -46 + -1x = 6 Add '46' to each side of the equation. -46 + 46 + -1x = 6 + 46 Combine like terms: -46 + 46 = 0 0 + -1x = 6 + 46 -1x = 6 + 46 Combine like terms: 6 + 46 = 52 -1x = 52 Divide each side by '-1'. x = -52 Simplifying x = -52